A university​ president's salary for the​ 2012-2013 academic year was $470,000 Calculate his​ 2013-2014 salary if he received a 2.7% raise.


1
Expert's answer
2022-01-16T07:44:06-0500

A university​ president's salary for the​ 2012-2013 academic year was $470,000 Calculate his​ 2013-2014 salary if he received a 2.7% raise.

Answer

Since 2013-2014 salary is a raise the percentage becomes 102.7%


Salary received in 2013-2014 =

470,000 × (102.7/100)=$482,690